### v1

@@ Appendix 0.A Pseudo Training Algorithm
Appendix 0.B Related Work
0.B.1 Diffusion-based Video Editing
@@
There has been considerable progress in adapting the achievements of diffusion-based image editing for video generative tasks. Compared to text-conditioned image generation, creating videos based solely on text introduces the complex challenge of producing temporally consistent and natural motion. In the absence of publicly available text-to-video diffusion models, Tune-A-Video [38] was at the forefront of one-shot based video editing. It proposes to inflate image diffusion model to pseudo video diffusion model by appending temporal modules to image diffusion model [28] and reprogramming spatial self-attention to spatio-temporal self-attention. Following this adaptation, the attention modules’ query projection matrices are fine-tuned on the input video. To eliminate the need for customizing model weights for every new video, various zero-shot editing methods have been developed. One approach involves guiding the generation process with attention maps, such as the injection of self-attention maps obtained from previous frames [5] or input video [25]. Another prevalent method integrates explicit structural cues, like depth or edge maps, into the reverse diffusion process. For instance, ControlNet [46] has been adapted for the video domain, facilitating the creation of structurally consistent frames in video generation [18] and translation tasks [15, 47]. Furthermore, GLIGEN’s [21] adaptation to the video domain by Ground-A-Video [17] demonstrates the use of both spatially-continuous depth map and spatially-discrete bounding box conditions, achieving multi-attribute editing of videos in a zero-shot manner.
@@
The advent of open-source text-to-video diffusion models [34, 33, 6, 35] has spurred research into separating, altering, and combining the appearance and motion elements of videos [49, 16, 36, 1, 41, 42, 27]. MotionDirector [49] and DreamVideo [36] have each suggested approaches for dividing fine-tuning processes into distinct learning phases for subject appearance and temporal motion, utilizing efficient fine-tuning methods. On the other hand, VMC [16] focuses on distilling the motion within a video by calculating the residual vectors between consecutive frames. In their work, they fine-tune temporal attention layers within cascaded video diffusion models to synchronize the ground-truth motion vector with the denoised motion vector estimate, successfully generating videos that replicate the motion pattern of an input video within diverse visual scenarios. Similarly, [42] introduces a space-time feature loss that constructs self-similarity matrices based on the differences in attention features between frames. This approach aims to minimize the discrepancy in self-similarity between the input and output videos.
0.B.2 Frequency-aware Visual Generation
@@
Spectral analysis plays a pivotal role in the domain of visual understanding and generation, offering insights into the temporal-spatial structure of pixel-domain frames through frequency-domain signals. [23] introduced a hierarchical motion estimation algorithm employing complex discrete wavelet transforms, effectively utilizing phase differences among subbands to indicate local translations within video frames. [29] enhanced scalable video compression through motion-compensated wavelet transforms, integrating a continuous deformable mesh motion model to achieve superior compression efficiency and motion representation.
@@
Furthermore, these spectral insights have been instrumental in refining algorithms and deepening architectural understanding, particularly within the contexts of U-Net and autoencoder frameworks. [43] provided a groundbreaking reinterpretation of deep learning architectures for image reconstruction, establishing a connection between deep learning and classical signal processing theories, including wavelets and compressed sensing. [44] developed a wavelet-based correction method, WCT2, to augment photorealism in style transfer, leveraging whitening and coloring transforms to preserve structural integrity and statistics within the VGG feature space.
@@
In the realm of U-Net in diffusion models, [37] highlighted the rapid dominance of noise over high-frequency information in residual U-Net denoisers. Concurrently, [19] proposed the Hybrid Video Diffusion Model (HVDM), a novel architecture that captures spatiotemporal dependencies using a disentangled representation combining 2D projection and 3D convolutions with wavelet decomposition. While [19] aims to reconstruct input video with frequency matching loss, our approach does not aim to reconstruct input, focusing instead on learning motion dynamics through motion vectors, thereby distinguishing our method within the landscape of spectral analysis applications in motion estimation and transfer.
Appendix 0.C Experimental details
@@
For spectral global motion alignment, we mainly use Haar wavelets with the number of levels $l=3$ for 8-frame videos and $l=4$ for 16-frame videos. We use DWT1DForward function from the PyTorch package [^1] https://pytorch-wavelets.readthedocs.io/en/latest/index.html. For spectral local motion refinement, we fix $\delta=0.05$ in frequency domain weighting $\omega(a,b)$. We set $\lambda_{g}=0.4$ and $\lambda_{l}=0.2$ for many cases, where we recommend to fine-tune $\lambda_{g}$ in a range of $[0.2,0.5]$, and $\lambda_{g}$ in a range of $[0.1,0.3]$. We follow other configurations, e.g. optimization algorithm, learning rate, training steps, etc., from the original motion transfer frameworks.
Appendix 0.D Additional Results
@@
This section provides additional qualitative comparisons of SMA across different baseline approaches. In Figure 8, we compare the efficacy of MotionDirector [49] in transferring motion, both with SMA integrated and without. Figures 9 and 10 delve into VMC’s [16] capabilities in customizing motion, showcasing outcomes with SMA and without its application. Additionally, Figure 11 contrasts the ability of Tune-A-Video and ControlVideo [38] to replicate the original motion, examining the impact of incorporating SMA. All our qualitative results can be viewed on our project page in video format.
Figure 8: Additional comparison within MotionDirector framework.
@@
Figure 11: Additional comparison within Tune-A-Video and ControlVideo frameworks.

### v2

@@ Appendix A Pseudo Training Algorithm
Appendix B Related Work
B.1 Diffusion-based Video Editing
@@
There has been considerable progress in adapting the achievements of diffusion-based image editing for video generative tasks. Compared to text-conditioned image generation, creating videos based solely on text introduces the complex challenge of producing temporally consistent and natural motion. In the absence of publicly available text-to-video diffusion models, Tune-A-Video (Wu et al. 2023) was at the forefront of one-shot based video editing. It proposes to inflate image diffusion model to pseudo video diffusion model by appending temporal modules to image diffusion model (Rombach et al. 2022) and reprogramming spatial self-attention to spatio-temporal self-attention. Following this adaptation, the attention modules’ query projection matrices are fine-tuned on the input video. To eliminate the need for customizing model weights for every new video, various zero-shot editing methods have been developed. One approach involves guiding the generation process with attention maps, such as the inekction of self-attention maps obtained from the input video (Ceylan, Huang, and Mitra 2023; Qi et al. 2023). Another prevalent method integrates explicit structural cues, like depth or edge maps, into the reverse diffusion process. For instance, ControlNet (Zhang, Rao, and Agrawala 2023) has been adapted for the video domain, facilitating the creation of structurally consistent frames in video generation (Khachatryan et al. 2023) and translation tasks (Hu and Xu 2023; Zhang et al. 2023b). Furthermore, GLIGEN’s (Li et al. 2023) adaptation to the video domain by Ground-A-Video (Jeong and Ye 2023) demonstrates the use of both spatially-continuous depth map and spatially-discrete bounding box conditions, achieving multi-attribute editing of videos in a zero-shot manner.
@@
The advent of open-source text-to-video diffusion models (Wang et al. 2023a; Sterling 2023; Chen et al. 2023a; Wang et al. 2023b) has spurred research into separating, altering, and combining the appearance and motion elements of videos (Zhao et al. 2023b; Jeong, Park, and Ye 2023; Wei et al. 2023; Bai et al. 2024; Yang et al. 2024; Yatim et al. 2023; Ren et al. 2024; Jeong et al. 2024). MotionDirector (Zhao et al. 2023b) and DreamVideo (Wei et al. 2023) have each suggested approaches for dividing fine-tuning processes into distinct learning phases for subject appearance and temporal motion, utilizing efficient fine-tuning methods. On the other hand, VMC (Jeong, Park, and Ye 2023) focuses on distilling the motion within a video by calculating the residual vectors between consecutive frames. In their work, they fine-tune temporal attention layers within cascaded video diffusion models to synchronize the ground-truth motion vector with the denoised motion vector estimate, successfully generating videos that replicate the motion pattern of an input video within diverse visual scenarios. Similarly, (Yatim et al. 2023) introduces a space-time feature loss that constructs self-similarity matrices based on the differences in attention features between frames. This approach aims to minimize the discrepancy in self-similarity between the input and output videos.
B.2 Frequency-aware Visual Generation
@@
Spectral analysis plays a pivotal role in the domain of visual understanding and generation, offering insights into the temporal-spatial structure of pixel-domain frames through frequency-domain signals. (Magarey and Kingsbury 1998) introduced a hierarchical motion estimation algorithm employing complex discrete wavelet transforms, effectively utilizing phase differences among subbands to indicate local translations within video frames. (Secker and Taubman 2002) enhanced scalable video compression through motion-compensated wavelet transforms, integrating a continuous deformable mesh motion model to achieve superior compression efficiency and motion representation.
@@
Furthermore, these spectral insights have been instrumental in refining algorithms and deepening architectural understanding, particularly within the contexts of U-Net and autoencoder frameworks. (Ye, Han, and Cha 2018) provided a groundbreaking reinterpretation of deep learning architectures for image reconstruction, establishing a connection between deep learning and classical signal processing theories, including wavelets and compressed sensing. (Yoo et al. 2019) developed a wavelet-based correction method, WCT2, to augment photorealism in style transfer, leveraging whitening and coloring transforms to preserve structural integrity and statistics within the VGG feature space.
@@
In the realm of U-Net in diffusion models, (Williams et al. 2024) highlighted the rapid dominance of noise over high-frequency information in residual U-Net denoisers. Concurrently, (Kim et al. 2024) proposed the Hybrid Video Diffusion Model (HVDM), a novel architecture that captures spatiotemporal dependencies using a disentangled representation combining 2D projection and 3D convolutions with wavelet decomposition. While (Kim et al. 2024) aims to reconstruct input video with frequency matching loss, our approach does not aim to reconstruct input, focusing instead on learning motion dynamics through motion vectors, thereby distinguishing our method within the landscape of spectral analysis applications in motion estimation and transfer.
Appendix C Experimental details
@@
For spectral global motion alignment, we mainly use Haar wavelets with the number of levels $l=3$ for 8-frame videos and $l=4$ for 16-frame videos. We use DWT1DForward function from the PyTorch package [^1] https://pytorch-wavelets.readthedocs.io/en/latest/index.html. For spectral local motion refinement, we fix $\delta=0.05$ in frequency domain weighting $\omega(a,b)$. We set $\lambda_{g}=0.4$ and $\lambda_{l}=0.2$ for many cases, where we recommend to fine-tune $\lambda_{g}$ in a range of $[0.2,0.5]$, and $\lambda_{g}$ in a range of $[0.1,0.3]$. We follow other configurations, e.g. optimization algorithm, learning rate, training steps, etc., from the original motion transfer frameworks for a fair comparisons. Code is included as a supplementary material.
Appendix D Additional analysis
Ablation study.
We first provide additional qualitative ablation results of global alignment in Fig. 7 (baseline: VMC). Combined with quantiative ablation studies in the main paper, these demonstrates the effectiveness of both global and local motion alignment. Please refer to the Figure 6 in the main paper for more analysis on the local term.
Training progress.
To analyze the training progress further, we visualize the intermediate customization results with and without proposed spectral motion alignment. Notably, Fig. 8a illustrates the gradual improvements of spatial layout attributed to the proposed global/local motion alignment. Specifically, our layout progressively aligns better as training proceeds. Both global and local terms contribute to this alignment as both match motion vectors $\delta\boldsymbol{v}_{0}$ encoding structural movement information.
Alternative frequency-domain approaches.
Our framework highlights two key insights: (a) susceptibility of 2D motion information in $\delta\boldsymbol{v}_{0}$ to spatial/motion artifacts, and (b) potential benefits of multi-scale decomposition for refining motion representation. In this context, both multi-scale Fourier and wavelet analyses offer theoretically viable options for spectral motion refinement. Demonstrating this, we use 2D Discrete Wavelet Transform (DWT) with Daubechies 3 wavelet, to obtain wavelet coefficients of ground truth and predicted motion vectors. Then each motion vector is reconstructed from these coefficients, excluding the finest high-frequency detail coefficients. Figure 8b illustrates that DWT-based refinement also significantly improves motion transfer compared to baseline (VMC), particularly by prioritizing low-frequency components of local motion encoding core motion information. While both methods show notable improvements, DFT outperforms DWT, with DWT requiring more fine-tuning of hyperparameters (e.g., the number of levels depending on spatial resolution, wavelet type, etc). Thus, we opt for DFT in refinement.
Appendix E Additional Results
@@
This section provides additional qualitative comparisons of SMA across different baseline approaches. Figures 9 and 10, 11 delve into VMC’s (Jeong, Park, and Ye 2023) capabilities in customizing motion, showcasing outcomes with SMA and without its application. In Figure 12, we compare the efficacy of MotionDirector (Zhao et al. 2023b) in transferring motion, both with SMA integrated and without. Additionally, Figure 13 contrasts the ability of ControlVideo (Wu et al. 2023) to replicate the original motion, examining the impact of incorporating SMA.
Figure 7: Ablation study on global alignment. Global motion alignment facilitates motion transfer.
Figure 8: (a) Visualization of the training progress with and without local term (Global motion alignment leads to similar trends). (b) Comparisons of local motion refinement with 2D DWT and 2D DFT (originally proposed).
@@
Figure 12: Additional comparison within MotionDirector method.
Figure 13: Additional comparison within ControlVideo frameworks.